OpenStudy (anonymous):

what is the solution to the following system of equations? y=x^2+x+8; y=x^2+2x+5

OpenStudy (anonymous):

You just have to use either linear combination or substitution. Hope you can get it from there.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

How do i do that?

OpenStudy (anonymous):

U could set them equal to each other like this... x^2+x+8 = x^2+2x+5 Solve for x. Then, plug in the value u get for x in any of the two equations to find y!

OpenStudy (anonymous):

x^2+x+8 = x^2+2x+5 solve it you'll get x = 3 and then put this value of "x" in any of equation.. you'll get y=20

OpenStudy (anonymous):